# On the Power of the Weisfeiler-Leman Test for Graph Motif Parameters

### v1

@@ Appendix B Technical Details for Section 3
The goal of this section is to show that for any labeled graph $F$ that has treewidth greater than $k$, there are labeled graphs $G$ and $H$ such that $G\equiv_{kWL}H$ but $\mathsf{homs}(F,G)\neq\mathsf{homs}(F,H)$. Thus $G$ and $H$ are witnesses to the fact that the $k$WL test cannot express the function $\mathsf{homs}(F,\cdot)$.
To do so, we show that two important results on the functions $\mathsf{homs}$ for unlabeled graphs still hold in the labeled setting. To distinguish more clearly between the labeled and unlabeled case we will use the term plain graphs to be those labeled graphs where the vertex- and edge-label alphabets are singletons. For technical reasons we we will initially focus on edge-labeled graphs, which are labeled graphs where the vertex-label alphabet is a singleton. For labeled graph $G$, we write $\mathsf{plain}(G)$ for the plain graph obtained by making all edge and vertex labels of $G$ the same (without changing the vertices or edges of $G$).
The main property of $\mathsf{homs}$ that we are interested in is equivalence for fixed classes of graphs. Formally, let $\mathcal{F}$ be a class of labeled graphs. We write $G\equiv_{\mathcal{F}}H$ if for every $F\in\mathcal{F}$ it holds that $\mathsf{homs}(F,G)=homs(F,H)$. Two families of graph classes will be particularly important in the following. We write $\mathcal{T}_{k}$ for the class of plain graphs with treewidth at most $k$, let $\mathcal{LT}_{k}$ be the class of labeled graphs of treewidth at most $k$, and let $\mathcal{ET}_{k}$ for the class of edge-labeled graphs with treewidth at most $k$.
@@
Our second key ingredient will be to show the same holds in the labeled setting. That is, that $\mathcal{LT}_{k}$ is labeled homomorphism-distinguishing closed. To simplify the most technical parts of the argument we in fact first show that $\mathcal{ET}_{k}$ is edge-labeled homomorphism-distinguishing closed.
@@
For plain graphs, Roberson showed that any class of graphs that is closed under weak oddomorphisms, disjoint unions, and restriction to connected components is homomorphism-distinguishing closed. It is one main goal of this section to show that the same also holds for edge-labeled graphs.
Theorem 15 (Edge-labeled Version of Theorem 6.2 (Roberson, 2022)).
Let $\mathcal{F}$ be a class of edge-labeled graphs such that
@@
Then $\mathcal{F}$ is edge-labeled homomorphism-distinguishing closed.
Before we move on to the proof of Theorem 15, we first observe that the class $\mathcal{ET}_{k}$ satisfies the conditions of the theorem. Indeed, Condition 15 is trivially satisfied as the treewidth of a graph is the maximum treewidth over the graph’s connected components (and labels do not influence treewidth). For Condition 1 we can reuse the following key part of the proof of Theorem 14.
@@
Recall that labels serve no explicit role in the definition of an oddomorphism beyond restricting the set of homomorphisms overall. Formalizing this observation we can show the same also for edge-labeled graphs.
@@
Let $F\in\mathcal{ET}_{k}$ and $G$ be a edge-labeled graph. If there is a weak oddomorphism from $F$ to $G$, then $G\in\mathcal{ET}_{k}$.
@@
First, observe that for labeled graphs $F,G$, if $\phi$ is an oddomorphism from $F$ to $G$, then $\phi$ is also an oddomorphism from $\mathsf{plain}(F)$ to $\mathsf{plain}(G)$. Now, if there is only a weak oddomorphism $\phi$ from $F$ to $G$, there is some subgraph $F^{\prime}$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$. Hence it is also a oddomorphism from $\mathsf{plain}(F^{\prime})$ to $\mathsf{plain}(G)$, meaning that where is a weak oddomorphism from $\mathsf{plain}(F)$ to $\mathsf{plain}(G)$.
We have that $tw(F)=tw(\mathsf{plain}(F))\leq k$, and therefore by Proposition 16 $\mathsf{plain}(G)\in\mathcal{T}_{k}$. And therefore also $tw(G)=tw(\mathsf{plain}(G))\leq k$, i.e., $G\in\mathcal{ET}_{k}$. ∎
@@
$\mathcal{ET}_{k}$ is edge-labeled homomorphism distinguishing closed.
The rest of this section is organized as follows. We first verify that Theorem 15 holds in Section B.1. We then argue Lemma 5 in Section B.3. Finally, we lift Theorem 18 to the labeled homomorphism distinguishing closedness of $\mathcal{LT}_{k}$ in Section B.2, thus proving Lemma 3. Finally, in Section B.4 we briefly note why Lemma 4 also holds in the labeled setting.
@@ B.1 Proof of Theorem 15
We will closely follow the proof of Theorem 6.2 in (Roberson, 2022). We will see that adapting the initial definitions in the right way will actually leave most of the framework for plain graphs intact. In many situations the effect of the edge-labels is on the argument is very subtle. For the sake of verifiability we therefore repeat the (slightly, in parts) modified arguments for the critical path in the unlabeled setting here. We wish to emphasize that, if not stated otherwise, all proofs are effectively due to Roberson.
@@
The graph has edge $\{(v,S),(u,T)\}$ in $E(\mathsf{CFI}(G,U))$ if and only if $\{v,u\}\in E(G)$ and $\{v,u\}\not\in S\Delta T$.
The definition can be naturally adapted to the edge-labeled setting by simply inheriting the labels of the original edges in $G$. To this end, let $G$ now be a edge-labeled graph. We will define the labeled graph $\mathsf{CCFI}(G,U)$ with $V(\mathsf{CCFI}(G,U))=V(\mathsf{CFI}(G,U))$, that is, the vertices are unaffected by the labels. For the edge-labeling function $\kappa$ edges we simply say that $\kappa(\{(v,S),(u,T)\})=\kappa(\{v,u\})$. That is, $\kappa(\{(v,S),(u,T)\})=c$ if and only if $\kappa(\{v,u\})=c$ and $\{v,u\}\not\in S\Delta T$.
We will show that this definition preserves the key properties of $\mathsf{CFI}$ graphs in the edge-labeled setting. We will say that two vertices $u,v$ are adjacent via label $i$ if there is an edge $\{u,v\}$ with label $i$. In fact, this will require almost no technical changes.
@@
As in the plain graph case, it is straightforward to verify that this is a bijection. Suppose that $(a,S)$ and $(b,T)$ are adjacent via label $i$ in $\mathsf{CCFI}(G,U)$, i.e., $\{a,b\}\in E_{i}(G)$ and $\{a,b\}\not\in S\Delta T$. Let $S^{\prime},T^{\prime}$ such that $\phi((a,S))=(a,S^{\prime})$, $\phi((b,T))=(b,T^{\prime})$. If $\{a,b\}=e$, then $S^{\prime}=S\Delta\{e\}$ and $T^{\prime}=T\Delta\{e\}$, hence $S\Delta T^{\prime}=S\Delta T$. Thus, $\{a,b\}$ is also not in $S^{\prime}\Delta T^{\prime}$ and $\phi((a,S))$ is adjacent to $\phi((b,T))$. In the other case that $\{a,b\}\neq e$, then $S^{\prime}$ and $T^{\prime}$ will both, individually, contain $\{a,b\}$ exactly if $S$ and $T$, respectively, did. So again $\{a,b\}\not\in S^{\prime}\Delta T^{\prime}$. In both cases, the labels of the edge in $\mathsf{CCFI}(G,U)$ and $\mathsf{CCFI}(G,U^{\prime})$ is simply inherited from $\{a,b\}$. Thus $\phi$ preserves labeled adjacency.
@@
As a next step we adapt (Roberson, 2022, Lemma 3.4) to the edge-labeled setting. We first recall a definition from Roberson (2022) that carries over unchanged from the plain graph setting. For any $U\subseteq V(G)$, the mapping $(v,S)\mapsto v$ is a homomorphism from $\mathsf{CCFI}(G,U)$ to $G$. We will refer to this mapping as $\rho$ in the following. For a homomorphism $\phi\in\mathsf{Hom}(F,G)$ define
$\mathsf{Hom}_{\phi}(F,\mathsf{CCFI}(G,U))=\{\psi\in\mathsf{Hom}(F,\mathsf{CCFI}(G,U)\mid\rho\circ\psi=\phi\}$
and observe that the sets $Hom_{\phi}(F,\mathsf{CCFI}(G,U)$ for $\phi\in\mathsf{Hom}(F,G)$ partition $\mathsf{Hom}(F,\mathsf{CCFI}(G,U))$.
The following statement is subtle in our context. Except for qualifying graphs $G$ and $F$ to be edge-labeled, the statement is verbatim the same as Lemma 3.4 in (Roberson, 2022). That is, the labels are not explicitly considered in the system of equations. Note however that the equations are relative to some homomorphism from $F$ to $G$, which makes the second set of equations implicitly follow the labeling. The fact that this system of equations does not change is the key to requiring only few changes in the later lemmas, as the argument proceeds mainly algebraically via this system of equations.
Lemma 20 (Edge-labeled version of Lemma 3.4 (Roberson, 2022)).
Let $G$ be a connected edge-labeled graph, let $U\subseteq V(G)$, and let $F$ be any edge-labeled graph. For a given $\phi\in\mathsf{Hom}(F,G)$, define variables $x_{e}^{a}$ for all $a\in V(F)$ and $e\in I(\phi(a)$. Then the elements of $Hom_{\phi}(F,\mathsf{CCFI}(G,U)$ are in bijection with solutions of the following equations over $\mathbb{Z}_{2}$:
@@
We first show that $\psi\in\mathsf{Hom}_{\phi}(F,\mathsf{CCFI}(G,U))$. By Equation 2, $\psi(a)$ is guaranteed to be a vertex of $\mathsf{CCFI}(G,U)$ for all $a\in V(F)$. To see that $\psi$ preserves labeled adjacency, suppose $\{a,b\}\in E_{i}(F)$. Then $e=\{\phi(a),\phi(b)\}\in E_{i}(G)$ since $\phi\in\mathsf{Hom}(F,G)$. From Equation 3 it follows that $x_{e}^{a}=x_{e}^{b}$. Then, by construction of $\mathsf{CCFI}(G,U)$, $\psi(a)$ is adjacent via label $i$ to $\psi(b)$ exactly if $\{\phi(a),\phi(b)\}\not\in S(a)\Delta S(b)$. Meaning that $\psi(a)$ is adjacent via label $i$ to $\psi(b)$ if $e$ is either in both $S(a)$ and $S(b)$ or neither, which is true because $x_{e}^{a}=x_{e}^{b}$.
@@
Since we can use the same system of equations (which does not explicitly consider the edge labels) as used in Roberson (2022), the development from Section 3 of (Roberson, 2022) holds almost unchanged from here. For the sake of completeness we repeat the relevant parts for the proof of Theorem 15 here, including some discussion on why the proofs are unaffected by the labels.
Let $G$ be a connected edge-labeled graph, $F$ a edge-labeled graph, and let $\phi\in\mathsf{Hom}(F,G)$. Let $R$ be the set of pairs $(a,e)$ such that $a\in V(F)$ and $e\in E(\phi(a))$. We define the matrices $A^{\phi}\in\mathbb{Z}_{2}^{V(F)\times R}$ and $B^{\phi}\in\mathbb{Z}_{2}^{E(F)\times R}$ as:
@@
Theorem 21 (Edge-labeled version of Theorem 3.6 (Roberson, 2022)).
Let $G$ be a connected edge-labeled graph, let $U\subseteq V(G)$, and let $\phi\in\mathsf{Hom}(F,G)$ for some edge-labeled graph $F$. Then $\mathsf{homs}_{\phi}(F,\mathsf{CCFI}(G))>0$ and
@@
Theorem 22 (Edge-labeled Version of Theorem 3.13 (Roberson, 2022)).
Let $G$ be a connected edge-labeled graph and $F$ be a edge-labeled graph. Then $\mathsf{homs}(F,\mathsf{CCFI}(G))\geq\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G))$. Furthermore, the inequality is strict if and only if there exists a weak oddomorphism from $F$ to $G$. If such a weak oddomorphism $\phi$ exists, there is a connected subgraph $F^{\prime}$ of $F$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$.
@@
The original proof of (Roberson, 2022, Theorem 3.13) works unchanged, although it might not be considered straightforward to observe that this is the case. We therefore repeat the key steps here with emphasis on the points where it needs to be observed that edge labels do not affect the argument. We wish to stress again that this proof is due to Roberson and we repeat it here only to allow for a more targeted discussion of why labels have no effect on the argument.
@@
Observe that a solution $(y,z)^{T}$ to this system satisfies $(A^{\phi})^{T}y=(B^{\phi})^{T}$. Let $O=\{a\in V(F)\mid y_{a}=1\}$ and let $F^{\prime}$ be the subgraph of $F$ with all vertices and edges $E(F^{\prime})=\{e\in E(F)\mid z_{e}=1\}$. The goal of the argument now is to show that $\phi$ is an oddomorphism from $F^{\prime}$ to $G$ (and thus a weak oddomorphism w.r.t. $F$). Since $F^{\prime}$ is a subgraph of $F$, it is clear that $\phi$ is still a homomorphism from $F^{\prime}$ to $G$, regardless of edge labels. The arguments for the properties of the oddomorphism themselves are parity arguments on the number of adjacent edges between $\phi^{-1}(v)$ and $\phi^{-1}(u)$ for adjacent $v,u\in V(G)$. These arguments also hold unchanged in the edge-labeled setting as the labeled homomorphism $\phi$ still preserves all adjacencies.Finally, $F^{\prime}$ as constructed above may not be connected. As in the unlabeled case we can then replace $F^{\prime}$ be a connected component by showing that if there is a weak oddomorphism $\phi|_{V(F^{\prime})}$ for some subgraph $F^{\prime}$, then we can assume w.l.o.g., that $F^{\prime}$ is connected (Roberson, 2022, Lemma 3.12). ∎
@@
Let $F,G$ be edge-labeled graphs. Then $\mathsf{homs}(F,\mathsf{CCFI}(G))>\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G))$.
@@
It is known that the identity $\mathsf{id}$ on $V(G)$ is an oddomorphism from $G$ to $G$ in the plain graph case (Roberson, 2022, Lemma3.14). The definition of oddomorphism does not take the labels of edges into account (only the neighborhoods of vertices, which are unaffected by labels), and $\mathsf{id}$ is still a homomorphism in the edge-labeled setting. Hence, $\mathsf{id}$ must also be an oddomorphism from $G$ to $G$ in the edge-labeled setting. The statement then follows immediately from Theorem 22. ∎
@@
This concludes our replication of the key statements of (Roberson, 2022, Section 3). We will require one additional statement that is not necessary in Roberson’s original argument for the proof of Theorem 15. For technical reasons we will require an edge-labeled graph $G$ such that for some $F$ we can guarantee that $\mathsf{homs}(F,G)>0$. For plain graphs this is straightforward by taking a large enough clique for $G$. In the edge-labeled case we need an alternative gadget.
@@
Let $\Delta$ be a finite edge-label alphabet and $n\geq 1$. There is a connected edge-labeled graph $G$ such that for every edge-labeled graph $F$ with at edge-label alphabet $\Delta$ and $n$ vertices, $\mathsf{homs}(F,G)>0$.
@@
Let $\mathcal{C}$ be the set of all (up to isomorphism) edge-labelings of the $n$ vertex clique with labels $\Delta$. Let $G^{*}$ be the disjoint union of all graphs in $\mathcal{C}$. To get the desired graph $G$, fix an arbitrary vertex of each component of $G^{*}$ and replace them all by a single vertex with the same adjacency relations. Any $n$ vertex graph $F$ with edge-labels in $\Delta$ is a subgraph of some element of $G^{\prime}\in\mathcal{C}$, and hence $\mathsf{homs}(F,G^{\prime})>0$. Since $G^{\prime}$ is a subgraph of $G$, also $\mathsf{homs}(F,G)>0$. ∎

### v3

@@ Appendix B Technical Details for Section 3
The goal of this section is to show that for any labeled graph $F$ that has treewidth greater than $k$, there are labeled graphs $G$ and $H$ such that $G\equiv_{kWL}H$ but $\mathsf{homs}(F,G)\neq\mathsf{homs}(F,H)$. Thus, $G$ and $H$ are witnesses to the fact that the $k$WL test cannot express the function $\mathsf{homs}(F,\cdot)$.
To do so, we show that two important results on the functions $\mathsf{homs}$ for unlabeled graphs still hold in the labeled setting. To distinguish more clearly between the labeled and unlabeled case we will use the term plain graphs to be those labeled graphs where the vertex- and edge-label alphabets are singletons. For technical reasons we will initially focus on edge-labeled graphs, which are labeled graphs where the vertex-label alphabet is a singleton. For labeled graph $G$, we write $\mathsf{plain}(G)$ for the plain graph obtained by making all edge and vertex labels of $G$ the same (without changing the vertices or edges of $G$).
The main property of $\mathsf{homs}$ that we are interested in is equivalence for fixed classes of graphs. Formally, let $\mathcal{F}$ be a class of labeled graphs. We write $G\equiv_{\mathcal{F}}H$ if for every $F\in\mathcal{F}$ it holds that $\mathsf{homs}(F,G)=\mathsf{homs}(F,H)$. Two families of graph classes will be particularly important in the following. We write $\mathcal{T}_{k}$ for the class of plain graphs with treewidth at most $k$, let $\mathcal{LT}_{k}$ be the class of labeled graphs of treewidth at most $k$, and let $\mathcal{ET}_{k}$ for the class of edge-labeled graphs with treewidth at most $k$.
@@
Our second key ingredient will be to show the same holds in the labeled setting. That is, that $\mathcal{LT}_{k}$ is labeled homomorphism-distinguishing closed.
@@
For plain graphs, Roberson showed that any class of graphs that is closed under weak oddomorphisms, disjoint unions, and restriction to connected components is homomorphism-distinguishing closed. It is one main goal of this section to show that the same also holds for labeled graphs.
Theorem 15 (Labeled Version of Theorem 6.2 (Roberson, 2022)).
Let $\mathcal{F}$ be a class of labeled graphs such that
@@
Then $\mathcal{F}$ is labeled homomorphism-distinguishing closed.
Before we move on to the proof of Theorem 15, we first observe that the class $\mathcal{LT}_{k}$ satisfies the conditions of the theorem. Indeed, Condition 15 is trivially satisfied as the treewidth of a graph is the maximum treewidth over the graph’s connected components (and labels do not influence treewidth). For Condition 1 we can reuse the following key part of the proof of Theorem 14.
@@
Recall that labels serve no explicit role in the definition of an oddomorphism beyond restricting the set of homomorphisms overall. Formalizing this observation we can show the same also for labeled graphs.
@@
Let $F\in\mathcal{LT}_{k}$ and $G$ be a labeled graph. If there is a weak oddomorphism from $F$ to $G$, then $G\in\mathcal{LT}_{k}$.
@@
First, observe that for labeled graphs $F,G$, if $\phi$ is an oddomorphism from $F$ to $G$, then $\phi$ is also an oddomorphism from $\mathsf{plain}(F)$ to $\mathsf{plain}(G)$. Now, if there is only a weak oddomorphism $\phi$ from $F$ to $G$, there is some subgraph $F^{\prime}$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$. Hence, it is also a oddomorphism from $\mathsf{plain}(F^{\prime})$ to $\mathsf{plain}(G)$, meaning that where is a weak oddomorphism from $\mathsf{plain}(F)$ to $\mathsf{plain}(G)$.
We have that $tw(F)=tw(\mathsf{plain}(F))\leq k$, and therefore by Proposition 16 $\mathsf{plain}(G)\in\mathcal{T}_{k}$. And therefore also $tw(G)=tw(\mathsf{plain}(G))\leq k$, i.e., $G\in\mathcal{LT}_{k}$. ∎
@@
$\mathcal{LT}_{k}$ is labeled homomorphism distinguishing closed.
The rest of this section is organized as follows. We first verify that Theorem 15 holds in Section B.1. We then argue Lemma 5 in Section B.2. Finally, in Section B.3 we briefly note why Lemma 4 also holds in the labeled setting.
@@ B.1 Proof of Theorem 15
We will closely follow the proof of Theorem 6.2 in (Roberson, 2022). We will see that adapting the initial definitions in the right way will actually leave most of the framework for plain graphs intact. In many situations the effect of the labels is on the argument is very subtle. For the sake of verifiability we therefore repeat the modified arguments for the critical path in the unlabeled setting here. We wish to emphasize that, if not stated otherwise, all proofs are effectively due to Roberson.
@@
The graph has edge $\{(v,S),(u,T)\}$ in $E(\mathsf{CFI}(G,U))$ if and only if $\{v,u\}\in E(G)$ and $v,u\not\in S\Delta T$.
The definition can be naturally adapted to the labeled setting by simply inheriting the labels of the original edges and vertices in $G$. To this end, let $G$ now be a labeled graph. We will define the labeled graph $\mathsf{CCFI}(G,U)$ with $V(\mathsf{CCFI}(G,U))=V(\mathsf{CFI}(G,U))$. For vertex $(v,S)\in V(G,U)$, we set the label $\lambda(v)$. For the edge-labeling function $\kappa$ edges we simply say that $\kappa(\{(v,S),(u,T)\})=\kappa(\{v,u\})$. That is, $\kappa(\{(v,S),(u,T)\})=c$ if and only if $\kappa(\{v,u\})=c$ and $v,u\not\in S\Delta T$.
We will show that this definition preserves the key properties of $\mathsf{CFI}$ graphs in the labeled setting. This might seem unintuitive as it ignores the labels on the sets $S\subseteq I(v)$ as well as the label of vertices in $U$. However, the key arguments are relative to homomorphisms on the graphs from which the $\mathsf{CCFI}$ construction is obtained. That is, these homomorphisms already follow the constraints imposed by the labelings. We will say that two vertices $u,v$ are adjacent via label $i$ if there is an edge $\{u,v\}$ with label $i$.
@@
As in the plain graph case, it is straightforward to verify that this is a bijection. Suppose that $(a,S)$ and $(b,T)$ are adjacent via label $i$ in $\mathsf{CCFI}(G,U)$, i.e., $\{a,b\}\in E_{i}(G)$ and $\{a,b\}\not\in S\Delta T$. Let $S^{\prime},T^{\prime}$ such that $\phi((a,S))=(a,S^{\prime})$, $\phi((b,T))=(b,T^{\prime})$. If $\{a,b\}=e$, then $S^{\prime}=S\Delta\{e\}$ and $T^{\prime}=T\Delta\{e\}$, hence $S\Delta T^{\prime}=S\Delta T$. Thus, $\{a,b\}$ is also not in $S^{\prime}\Delta T^{\prime}$ and $\phi((a,S))$ is adjacent to $\phi((b,T))$. In the other case that $\{a,b\}\neq e$, then $S^{\prime}$ and $T^{\prime}$ will both, individually, contain $\{a,b\}$ exactly if $S$ and $T$, respectively, did. So again $\{a,b\}\not\in S^{\prime}\Delta T^{\prime}$. In both cases, the labels of the edge in $\mathsf{CCFI}(G,U)$ and $\mathsf{CCFI}(G,U^{\prime})$ is simply inherited from $\{a,b\}$. Thus, $\phi$ preserves labeled adjacency.
@@
As a next step we adapt (Roberson, 2022, Lemma 3.4) to the labeled setting. We first recall a definition from Roberson (2022) that carries over unchanged from the plain graph setting. For any $U\subseteq V(G)$, the mapping $(v,S)\mapsto v$ is a homomorphism from $\mathsf{CCFI}(G,U)$ to $G$. We will refer to this mapping as $\rho$ in the following. For a homomorphism $\phi\in\mathsf{Hom}(F,G)$ define
$\mathsf{Hom}_{\phi}(F,\mathsf{CCFI}(G,U))=\{\psi\in\mathsf{Hom}(F,\mathsf{CCFI}(G,U))\mid\rho\circ\psi=\phi\}$
and observe that the sets $\mathsf{Hom}_{\phi}(F,\mathsf{CCFI}(G,U)$ for $\phi\in\mathsf{Hom}(F,G)$ partition $\mathsf{Hom}(F,\mathsf{CCFI}(G,U))$.
The following statement is subtle in our context. Except for qualifying graphs $G$ and $F$ to be labeled, the statement is verbatim the same as Lemma 3.4 in (Roberson, 2022). That is, the labels are not explicitly considered in the system of equations. Note however that the equations are relative to some homomorphism from $F$ to $G$, which makes the second set of equations implicitly follow the labeling. The fact that this system of equations does not change is the key to requiring only few changes in the later lemmas, as the argument proceeds mainly algebraically via this system of equations.
Lemma 20 (Labeled version of Lemma 3.4 (Roberson, 2022)).
Let $G$ be a connected labeled graph, let $U\subseteq V(G)$, and let $F$ be any labeled graph. For a given $\phi\in\mathsf{Hom}(F,G)$, define variables $x_{e}^{a}$ for all $a\in V(F)$ and $e\in I(\phi(a))$. Then the elements of $Hom_{\phi}(F,\mathsf{CCFI}(G,U))$ are in bijection with solutions of the following equations over $\mathbb{Z}_{2}$:
@@
We first show that $\psi\in\mathsf{Hom}_{\phi}(F,\mathsf{CCFI}(G,U))$. By Equation 2, $\psi(a)$ is guaranteed to be a vertex of $\mathsf{CCFI}(G,U)$ for all $a\in V(F)$. To see that $\psi$ preserves labeled adjacency, suppose $\{a,b\}\in E_{i}(F)$. Then $e=\{\phi(a),\phi(b)\}\in E_{i}(G)$ since $\phi\in\mathsf{Hom}(F,G)$. From Equation 3 it follows that $x_{e}^{a}=x_{e}^{b}$. Then, by construction of $\mathsf{CCFI}(G,U)$, $\psi(a)$ is adjacent via label $i$ to $\psi(b)$ exactly if $\{\phi(a),\phi(b)\}\not\in S(a)\Delta S(b)$. Meaning that $\psi(a)$ is adjacent via label $i$ to $\psi(b)$ if $e$ is either in both $S(a)$ and $S(b)$ or in neither, which is true because $x_{e}^{a}=x_{e}^{b}$. Vertex labels are directly preserved via $\phi$.
@@
Since we can use the same system of equations (which does not explicitly consider labels) as used in Roberson (2022), the development from Section 3 of (Roberson, 2022) holds almost unchanged from here. For the sake of completeness we repeat the relevant parts for the proof of Theorem 15 here, including discussion on why the proofs are unaffected by the labels where relevant.
Let $G$ be a connected labeled graph, $F$ a labeled graph, and let $\phi\in\mathsf{Hom}(F,G)$. Let $R$ be the set of pairs $(a,e)$ such that $a\in V(F)$ and $e\in E(\phi(a))$. We define the matrices $A^{\phi}\in\mathbb{Z}_{2}^{V(F)\times R}$ and $B^{\phi}\in\mathbb{Z}_{2}^{E(F)\times R}$ as:
@@
Theorem 21 (Labeled version of Theorem 3.6 (Roberson, 2022)).
Let $G$ be a connected labeled graph, let $U\subseteq V(G)$, and let $\phi\in\mathsf{Hom}(F,G)$ for some labeled graph $F$. Then $\mathsf{homs}_{\phi}(F,\mathsf{CCFI}(G))>0$ and
@@
Theorem 22 (Labeled Version of Theorem 3.13 (Roberson, 2022)).
Let $G$ be a connected labeled graph and $F$ be a labeled graph. Then $\mathsf{homs}(F,\mathsf{CCFI}(G))\geq\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G))$. Furthermore, the inequality is strict if and only if there exists a weak oddomorphism from $F$ to $G$. If such a weak oddomorphism $\phi$ exists, there is a connected subgraph $F^{\prime}$ of $F$ such that $\phi|_{V(F^{\prime})}$ is an oddomorphism from $F^{\prime}$ to $G$.
@@
The original proof of (Roberson, 2022, Theorem 3.13) works unchanged, although it might not be considered straightforward to observe that this is the case. We therefore repeat the key steps here with emphasis on the points where it needs to be observed that edge labels do not affect the argument. We wish to stress again that this proof is due to Roberson. We repeat it here to allow for a more targeted discussion of why labels have minimal effect on the argument, and to note necessary subtle differences.
@@
Observe that a solution $(y,z)^{T}$ to this system satisfies $(A^{\phi})^{T}y=(B^{\phi})^{T}$. Let $O=\{a\in V(F)\mid y_{a}=1\}$ and let $F^{\prime}$ be the subgraph of $F$ with all vertices and edges $E(F^{\prime})=\{e\in E(F)\mid z_{e}=1\}$. The goal of the argument now is to show that $\phi$ is an oddomorphism from $F^{\prime}$ to $G$ (and thus a weak oddomorphism w.r.t. $F$). Since $F^{\prime}$ is a subgraph of $F$, it is clear that $\phi$ is still a homomorphism from $F^{\prime}$ to $G$, regardless of labels. The arguments for the properties of the oddomorphism themselves are parity arguments on the number of adjacent edges between $\phi^{-1}(v)$ and $\phi^{-1}(u)$ for adjacent $v,u\in V(G)$. These arguments also hold unchanged in the labeled setting as the labeled homomorphism $\phi$ still preserves all adjacencies. Finally, $F^{\prime}$ as constructed above may not be connected. As in the unlabeled case we can then replace $F^{\prime}$ be a connected component by showing that if there is a weak oddomorphism $\phi|_{V(F^{\prime})}$ for some subgraph $F^{\prime}$, then we can assume w.l.o.g., that $F^{\prime}$ is connected (Roberson, 2022, Lemma 3.12). ∎
@@
Let $F,G$ be labeled graphs. Then $\mathsf{homs}(F,\mathsf{CCFI}(G))>\mathsf{homs}(F,\mathsf{CCFI}^{\times}(G))$.
@@
It is known that the identity $\mathsf{id}$ on $V(G)$ is an oddomorphism from $G$ to $G$ in the plain graph case (Roberson, 2022, Lemma3.14). The definition of oddomorphism does not take the labels of edges into account (only the neighborhoods of vertices, which are unaffected by labels), and $\mathsf{id}$ is still a homomorphism in the labeled setting. Hence, $\mathsf{id}$ must also be an oddomorphism from $G$ to $G$ in the labeled setting. The statement then follows immediately from Theorem 22. ∎
@@
This concludes our replication of the key statements of (Roberson, 2022, Section 3). We will require one additional statement that is not necessary in Roberson’s original argument for the proof of Theorem 15. For technical reasons we will require an labeled graph $G$ such that for some $F$ we can guarantee that $\mathsf{homs}(F,G)>0$. For plain graphs this is straightforward by taking a large enough clique for $G$. In the labeled case we need an alternative gadget.
@@
Let $\Sigma,\Delta$ be finite label alphabets and $n\geq 1$. There is a connected labeled graph $G$ such that for every labeled graph $F$ with at vertex/edge-label alphabets $\Sigma,\Delta$ and $n$ vertices, $\mathsf{homs}(F,G)>0$.
@@
Let $\mathcal{C}$ be the set of all (up to isomorphism) labelings of the $n$ vertex clique with labels from $\Sigma$ and $\Delta$. Let $G^{*}$ be the disjoint union of all graphs in $\mathcal{C}$. To get the desired graph $G$, add a fresh vertex with arbitrary label to $G^{*}$ that is adjacent to an arbitrary vertex of each connected component (via an arbitrary edge label). Any $n$ vertex graph $F$ with labels from $\Sigma,\Delta$ is a subgraph of some element of $G^{\prime}\in\mathcal{C}$, and hence $\mathsf{homs}(F,G^{\prime})>0$. Since $G^{\prime}$ is a subgraph of $G$, also $\mathsf{homs}(F,G)>0$. ∎